determine if the following series converge. if it converges, find its limit L (if possible)
5. Σ. In()
Answer #1

5. Ž inck) (by using Integral rest) The above series converges y and only I enx dx converges. ý 0 considera - det lim + l enxK-1 considel, an = nat! anti = n+! 0+1 = (D+1)2+1 D2 + 2n+2 um anti = neo an um nt no 12+2n+2 net n - eum lit the young (it k